For the gas constant, use R = 8.314 J/mol-K For A
Answer #1

Given data

Volume of the Syringe V = 47.5 cc = 4.75 * 10-5 m3

Pressure P = 1.58 atm = 1.58 *1.01 * 105 Pa = 1.59 *105Pa ( 1atm = 1.01 * 105 Pa)

Temperature T = 19.8 oC = (19.8 + 273) K =292.5K

R = 8.314 J/mol-K

Avogadro's number No = 6.022 * 1023 molecules

a) The number of moles in the Syringe

n = PV / RT

   =(1.59 * 105 * 4.75 * 10-5 ) / ( 8.314 * 292.5)

= 3.105 * 10-3 moles

Number of Molecules = Number of moles * Avogadro's number

   = 3.105 * 10-3 * 6.022 * 1023

   = 1.86 * 1021 molecules

Mass of the Air in the Syringe = Number of moles * Molar mass of air

   = 3.105*10-3 moles  * 28.97 *10-3 Kg/mol

   = 8.99 * 10-5 Kg

b) Temperature remains unchanged

Volume after compression Vf = 41.6 cc = 4.1 * 10-5 m3

Pf = ?

   PV = Pf Vf

Pf = PV / Vf

   = 1.59 * 105 * 4.75 * 10-5 / 4.16 * 10-5

= 1.81 * 105Pa

   = 1.81 * 105 / 1.01 *105

= 1.79 atm , 1 atm = 1.01*105 Pa